Q. Write the chemical equation of the two methods of obtaining ethanol with the help of the Griginard reagent, and the chemical equation of the two axes of making diethyl ether from ethyl O alcohol.
1
Expert's answer
2020-02-27T09:07:08-0500

Solution.

1)

"H-C(=O)-H + R-MgX + H2O = CH3CH2OH"

"CH2-(O)-CH2 + R-MgX + H2O = CH3CH2OH"

2)

"C2H5OH + H2SO4 = C2H5-O-SO3H + H2O"

"C2H5OH + C2H5-O-SO3H = C2H5-O-C2H5 + H2SO4"

Answer:

1)

"H-C(=O)-H + R-MgX + H2O = CH3CH2OH"

"CH2-(O)-CH2 + R-MgX + H2O = CH3CH2OH"

2)

"C2H5OH + H2SO4 = C2H5-O-SO3H + H2O"

"C2H5OH + C2H5-O-SO3H = C2H5-O-C2H5 + H2SO4"